OpenStudy (anonymous):

help please.. y = exp (√1 + 5x^3) get dy/dx

OpenStudy (anonymous):

is the radical over the whole thing? \[exp(\sqrt{1+5x^3})\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

very sloppy on my part

OpenStudy (anonymous):

derivative of exp(something) = exp(something) times derivative of something derivative of root(something) is one over two root (something)time derivative of something etc

OpenStudy (anonymous):

i dropped the 5 by mistake. should be \[\exp(\sqrt{1+5x^3})\times \frac{1}{2\sqrt{1+5x^3}}\times 15x^2\]